i have a problem with a question
x+y+z=4
X^2+y^2+z^2=14
x^3+y^3+z^3=34
x^4+y^4+z^4=......
answer it please

Answer by ankor@dixie-net.com(22740)

x+y+z=4
X^2+y^2+z^2=14
x^3+y^3+z^3=34
x^4+y^4+z^4=......
:
I came up with x = -1; y = 2; z = 3
:
Therefore
-1^4 + 2^4 + 3^4 =
+1 + 16 + 81 = 98
;
:
Check these values in the 2nd equation
-1^2 + 2^2 + 3^2 =
1 + 4 + 9 = 14
and
the 2nd equation
-1^3 + 2^3 + 3^3 =
-1 + 8 + 27 = 34